OpenStudy (anonymous):

I GIVE MEDALS :) What is the solution to the following equation? 5(2x − 6) + 20 = 10 9 5 3 2

TheSmartOne (thesmartone):

Hint: \(\sf a(b-c) =ab-ac\) So distribute \(\sf 5(2x-6)=?\)

TheSmartOne (thesmartone):

If you are getting confused with the 2x, then: another hint: \(\sf a(bx-c) = abx -ac\)